OpenStudy (anonymous):

Can anyone help me with the algebra word problem? A delivery truck is transporting boxes of two sizes: large and small. The combined weight of a large and small box is 80 pounds. The truck is transporting 60 large boxes and 55 small boxes. If the truck is carrying a total of 4625 pounds in boxes, how much does each type of box weigh? Large box: pound: ____ pounds Small box _____ pounds

OpenStudy (mertsj):

Let x be the weight of the small box. Then 80-x is the weight of the large box. Can you get it from there?

OpenStudy (anonymous):

No. This is a familiar question but this time I am searching for L the weight of the large box ans S the weight of the small box. I understand it is two equations but how to solve the correct one escapes me.

OpenStudy (anonymous):

60(80-x) + 55x = 4625 Can you get it from there

OpenStudy (anonymous):

If not I can help you with the problem

OpenStudy (anonymous):

Help!

OpenStudy (anonymous):

ok so I got 60(80-x) + 55x = 4625 4800 -60x + 55x = 4625 -60x + 55x = 4625 - 4800 -5x = -175 x = ?

OpenStudy (anonymous):

How's that

OpenStudy (anonymous):

I am calculating!

OpenStudy (anonymous):

So the answer would be that x = 35 Since the weight of the small boxes was x so the weight of the small boxes is 35 and the weight of the large boxes was (80-x) so the weight of the large boxes is 80 - 35 = 45 so the weight of the large boxes is 45 W of S = 35 pounds W of L = 45 pounds

OpenStudy (anonymous):

That is what I got but I could not tell which was S and which is L. How did you find the answer to that?

OpenStudy (anonymous):

well before we established that the weight of the small boxes was x and the weight of the large boxes was ( 80-x).

OpenStudy (anonymous):

We found out the value of x and plugged it in

OpenStudy (anonymous):

So there it is

OpenStudy (anonymous):

Dev you deserve a medal. Thank you. I will try my next word problem and see if I get it right. Thank you!

OpenStudy (anonymous):

Ok good luck and thanks

OpenStudy (anonymous):

No problem and thank you for showing me the steps.